# FusionMamba: Dynamic Feature Enhancement for Multimodal Image Fusion with Mamba

### v1

@@ Abstract.
Figure 1. Illustration of qualitative and quantitative results of multi-modal image fusion. (a) Qualitative visualization between classical U2Fusion (Guo et al., 2024) and our FusionMamba is shown in the second row while the sub-figures on the first row are source image pairs. (b) Quantitative results imply the efficacy of FusionMamba over other methods on various datasets and metrics.
@@ 1. Introduction
Different sensors capture unique data perspectives. Infrared sensors excel in thermal radiation capture, visible sensors in detailed texture data (Xie et al., 2023). Medical imaging’s MRI and CT provide structural insights (Zhou et al., 2023), while PET offers functional data (Tang et al., 2023b). Due to hardware limits, one sensor can’t capture the full image complexity (Liu et al., 2019). Image fusion integrates source images to create informative ones. Infrared-Visible and Medical Image Fusion are crucial sub-categories (Liu et al., 2019), aiding diagnosis and enhancing images (Ronneberger et al., 2015). Fusion addresses sensitivity and noise issues, supporting practical applications (Chen and Varshney, 2007; Yu et al., 2021; Ye et al., 2024; Yu et al., 2023).
@@
In recent years, there has been a significant increase in the use of deep learning for multimodal image fusion, primarily employing CNNs (Guo et al., 2024; Piella and Heijmans, 2003; Cvejic et al., 2005; Hwang et al., 2015) and Transformer architectures (Xie et al., 2023; Guan et al., 2023; Ma et al., 2022; Liu et al., 2024; Pei et al., 2024; Koroleva et al., 2005) for feature extraction and reconstruction. However, these models face limitations. However, both types of models have inherent limitations. Specifically, the utilization of static convolutional layers in these fusion methods results in uniform traversal of the entire image, limiting their ability to capture global contextual information due to constrained receptive fields (Ma et al., 2022). Moreover, the indiscriminate processing of each pixel by static convolutional layers overlooks nuanced spatial differences across various positions. Given the exigency of image fusion tasks necessitating diverse input data processing, current convolution-based techniques struggle to extract features effectively from different modalities, leading to diminished fusion performance. Transformer-based models excel in global modeling but suffer from quadratic complexity concerning image size due to their self-attention mechanism (Liu et al., 2024), imposing substantial computational overhead. Additionally, Transformers exhibit lower precision in capturing local features compared to CNNs. Although certain fusion models (Xie et al., 2023; Guan et al., 2023) adopt a hybrid approach by amalgamating convolutional and Transformer layers to leverage their respective strengths and mitigate weaknesses, computational overhead remains a significant concern.
@@
Image fusion is an essential image enhancement technique (Ma et al., 2022). Methodologies often employ three main feature fusion strategies. Firstly, some methods use element-wise operations like addition, multiplication, or concatenation (Xu et al., 2020b; Hwang et al., 2015) to merge features from different modalities. However, these approaches overlook inter-modal relationships, leading to compromised fusion performance. Secondly, there are approaches that focus on deeper feature extraction (Xie et al., 2023; Guan et al., 2023), but they still lack effective inter-modal interaction and texture detail emphasis. Lastly, specific techniques utilize cross-modal fusion (Ma et al., 2022), including convolution-based and Transformer-based methods, which partially address feature interaction but have their limitations. Overall, contemporary fusion methodologies struggle to optimally utilize modal features and highlight crucial information, indicating a need for improved modal connections and emphasis on key image details.
@@
The advancement of Mamba (Liu et al., 2024) represents a promising avenue for achieving a balance between global receptive fields and computational efficiency. By formalizing the discretized state-space equations of Mamba into recursive form and incorporating specially designed structured reparameterization (Ruan and Xiang, 2024), it becomes capable of simulating very long-range dependencies. Moreover, the implementation of a parallel scanning algorithm (Pei et al., 2024) enables Mamba to process each token in a parallel manner, thereby facilitating efficient training on modern hardware like GPUs (Guo et al., 2024). These encouraging characteristics inspire us to further explore Mamba’s potential for efficient long-range modeling within image fusion networks.
@@
For the feature fusion part, we design a Dynamic Feature Fusion Module comprising a Dynamic Feature Enhancement Module and a Cross Modality Fusion Mamba Module. Unlike previous feature fusion methods, this fusion module dynamically enhances the detail texture information and differential information in source images and facilitates better information interaction between modalities. The Dynamic Feature Enhancement Module (DFEM), composed of dynamic differential convolution and dynamic difference-sensing attention, is used for adaptive feature enhancement. The DFEM module dynamically enhances critical information by establishing connections between input features from different modalities. We design a Cross Modality Fusion Mamba Module to effectively mine correlation information between modalities. Our network architecture adopts a Unet (Zhao et al., 2023a) multi-layer structure, realizing an efficient and universal image fusion framework. Experimental results demonstrate that our proposed method outperforms state-of-the-art (SOTA) image fusion methods across various evaluation metrics on multiple multimodal baseline datasets, including IR-VIS fusion, CT-MRI fusion, PET-MRI fusion, SPECT-MRI fusion, and GFP-PC fusion.
@@ 2.1. Deep Multimodal Image Fusion
Image fusion, as a critical image enhancement technology, has garnered significant attention in recent years (Ma et al., 2022). In the era of deep learning, the primary image fusion methodologies can be categorized into four groups: Convolutional Neural Network (CNN) models (Guo et al., 2024; Piella and Heijmans, 2003; Cvejic et al., 2005; Chen and Varshney, 2007; Hwang et al., 2015), Generative Adversarial Network (GAN) models (Li et al., 2020, 2021), Autoencoder (AE) based models (Ram Prabhakar et al., 2017; Xu et al., 2020a; Zhang and Ma, 2021; Tang et al., 2022; Li et al., 2020; Li and Wu, 2018), and models based on Transformer architectures (Koroleva et al., 2005; Ma et al., 2020; Guan et al., 2023; Liu et al., 2024; Pei et al., 2024). These methods often employ straightforward fusion rules like element addition, averaging, or multiplication (Tang et al., 2022). While these convolutional methods can achieve satisfactory fusion results, they are limited in their ability to extract global features due to inherent convolution constraints (Ma et al., 2020). Ma et al. (Ma et al., 2019) introduced an end-to-end GAN model named FusionGAN, where the source image is fed into the generation network, and the fusion result is generated by the discriminator network. Subsequently, they proposed a dual discriminator conditional generative adversarial network (DDcGAN) (Tang et al., 2023b), which employs two discriminators to discern structural differences between the fused image and the two source images. While GANs demonstrate strong performance, their training process can be unstable, leading to potential texture distortion and difficulties in achieving stable fusion results due to the need to balance multi-source images using dual discriminators. Models based on autoencoders (AE) require a fusion rule to integrate features extracted from the encoder and generate the fusion result from the decoder. For example, RFN-Nest (Zhang and Ma, 2021) introduced a two-stage learning strategy and a customizable fusion approach. Furthermore, researchers have explored Transformer-based methods, with IFT (Liu et al., 2024) and CDDFuse (Tang et al., 2023a) effectively merging CNNs with Transformer architectures. SwinFusion (Ma et al., 2020) leverages SwinTransformer’s unique sliding window attention mechanism to address limitations in the convolutional fusion approach.
@@ 2.2. Mamba
State Space Models (SSMs) (Li and Wu, 2018) have become a competitive backbone in deep learning, originating from classic control theory and offering linear scalability with sequence length for long-range dependency modeling. Structured State Space Sequence Models (S4) and Mamba (Tang et al., 2023a), both rely on a classical continuous system that maps a 1D input function or sequence, denoted as $x(t)\in\mathbb{R}^{L}$, through intermediate implicit states $h(t)\in\mathbb{R}^{L}$ to an output $y(t)\in\mathbb{R}^{L}$. SSMs can be represented as the following linear Ordinary Differential Equation (ODE):
@@
Mamba has significantly advanced natural language tasks, surpassing traditional Transformers, with its data-dependent mechanisms, efficient hardware, and superior language processing. Expanding beyond language tasks, Mamba has also been applied successfully to vision tasks like image classification, video understanding, and biomedical image segmentation. This success has spurred a wave of research focusing on adapting Mamba-based models for specialized vision applications, including medical image segmentation with adaptations like Vm-unet (Tang et al., 2023a). Additionally, Mamba has been integrated into graph representation tasks, enhancing graph embeddings and processing capabilities through models. Mamba’s versatility and efficiency make it a compelling choice for a wide range of applications, from language processing to computer vision representation tasks (Tang et al., 2023b).
@@ 3.1. Motivation
The motivation of this paper lies in addressing the common problems and challenges in the field of multimodal image fusion. In recent years, with the recognition of the powerful capabilities of deep learning methods in image feature extraction, many fusion algorithms have emerged. However, the current mainstream convolutional neural networks (CNNs) (Piella and Heijmans, 2003; Cvejic et al., 2005; Chen and Varshney, 2007; Hwang et al., 2015) and Transformers models (Koroleva et al., 2005; Ma et al., 2020; Liu et al., 2024; Guan et al., 2023; Pei et al., 2024) have inherent limitations in image fusion, prompting us to delve into and design a new Dynamic Feature Enhancement Model to tackle these challenges and issues.
On the one hand, current fusion methods mainly use static convolutional layers for feature extraction and reconstruction, which have limitations in capturing nuanced spatial differences and global contextual information (Ma et al., 2020). On the other hand, Transformer models excel in global modeling, but their quadratic complexity in image size due to the self-attention mechanism restricts their computational efficiency (Tang et al., 2023a). Additionally, they are less accurate in capturing local features compared to CNNs. Moreover, existing fusion methods also fall short in feature fusion, unable to effectively extract features from different modalities, leading to a decline in fusion performance.
@@ 3.3. Dynamic Vision State Space Module
We propose the Dynamic Vision State Space (DVSS) module as a modification of the SSM block for image fusion processing. In Fig. 2(b), starting with the input deep feature, we initially apply LayerNorm (LN) followed by the Efficient State Space Module (ESSM) (Li and Wu, 2018) to capture spatial long-term dependencies.
Since SSMs process flattened feature maps as 1D token sequences, the chosen flattening strategy significantly impacts the number of neighboring pixels in the sequence. For example, when using the four-direction unfolding strategy, only four nearest neighbors are accessible to the anchor pixel. Specially, efficient 2D scan (ES2D, in Fig. 2(e)) (Li and Wu, 2018) efficiently downscales 2D-Selective-Scan (SS2D) using skipping sampling and combines processed patches for global feature extraction. Consequently, some spatially close pixels in the 2D feature map become distant from each other in the 1D token sequence, potentially resulting in local pixel forgetting. To tackle this issue, we introduce an additional dynamic local convolution after ES2D (Li and Wu, 2018) to restore neighborhood similarity. The Learnable Descriptive Convolution (LDC, in Fig. 2(f)) effectively learns intricate textural features, making it highly suitable in here. we employ LayerNorm for normalization and then utilize Learnable Descriptive Convolution (LDC) (Liu et al., 2019) to compensate for local features.
Moreover, SSMs often introduce a larger number of hidden states to remember very long-range dependencies, leading to significant channel redundancy when visualizing the activation results for different channels (Tang et al., 2023b). To enhance the expressive power of different channels, we integrate Efficient Channel Attention (ECA, in Fig. 2(g)) (Ruan and Xiang, 2024) into DVSS. This allows SSMs to concentrate on learning diverse channel representations, with subsequent channel attention selecting critical channels to prevent redundancy. DVSS is formulated as:
@@ 3.4. Dynamic Feature Enhancement Module
Within the module, the Learnable Descriptive Convolution (LDC)(Liu et al., 2019) module enhances texture processing on input feature maps using learnable mask parameters and convolution operations. By adjusting the convolution kernel’s weights, it emphasizes texture information, enhancing the model’s texture feature perception. Conversely, the Dynamic Difference Perception Attention calculates difference weights between input feature maps and applies them to both fused and original features, amplifying feature differences. This method helps the model effectively capture subtle differences among input features, leading to improved resolution and perception across different features and contributing to the model’s enhanced fusion performance. Finally, $\mathrm{DFEM}_{1}$ combines the texture enhancement feature, difference enhancement feature, and low-frequency feature ($\textbf{F}_{1}^{n}$) with background information, outputting $\textbf{D}_{1}^{n}$.
@@ 3.5. Cross Modality Fusion Mamba Module
These hybrid enhanced features are then input into the Efficient Spatial Scanning 2D (ES2D) (Li and Wu, 2018) layer to capture the spatial long-term dependency.
@@ 3.6. Loss Function
SSIM can measure the degree of distortion and the degree of similarity between two images (Chen and Varshney, 2007). We use SSIM loss to guarantee structure similarity between the fused image and source images. It can be defined as
@@ 4.1.1. Datasets
we choose the KAIST (Zhao et al., 2023b) dataset to train our FusionMamba. In this dataset, 70000 pairs of infrared and visible images are chosen for training. These images are also converted to gray scale and resized to 256 × 256. We use Harvard Medical Dataset  [^1] http://www.med.harvard.edu/AANLIB/home.html as the training and test dataset independently. There are 166 CT-MRI image pairs, 329 PET-MRI image pairs and 539 SPECT-MRI image pairs. The size of images is 256×256. We augment image pairs to 30000 via image rotation, which can improve the model and help to combat potential overfitting. 21 pairs of test images are randomly chosen to evaluate the model. The GFP database1 released by John Innes Centre (Zhang et al., 2020) is employed in our experiments. The database contains 148 pairs of preregistered GFP and PC images of Arabidopsis thealiana cells with the same size of 358 × 358 pixels. We also augment image pairs to 30000 via image rotation. 20 pairs of test images are randomly chosen to evaluate the model.

### v2

@@ Abstract.
Figure 1. Illustration of qualitative and quantitative results of multi-modal image fusion. (a) Qualitative visualization between classical U2Fusion (Xu et al., 2020a) and our FusionMamba is shown in the second row while the sub-figures on the first row are source image pairs. (b) Quantitative results imply the efficacy of FusionMamba over other methods on various datasets and metrics.
@@ 1. Introduction
Different sensors capture unique data perspectives. Infrared sensors excel in thermal radiation capture, visible sensors in detailed texture data (Zhang and Demiris, 2023). Medical imaging’s MRI and CT provide structural insights (Zhou et al., 2023), while PET offers functional data (3). Due to hardware limits, one sensor can’t capture the full image complexity (Liu et al., 2019). Image fusion integrates source images to create informative ones. Infrared-Visible and Medical Image Fusion are crucial sub-categories (Tang et al., 2023b), aiding diagnosis and enhancing images (Zhang and Demiris, 2023). Fusion addresses sensitivity and noise issues, supporting practical applications (Yu et al., 2021; Ye et al., 2024; Yu et al., 2023).
@@
In recent years, there has been a significant increase in the use of deep learning for multimodal image fusion, primarily employing CNNs (Xu et al., 2020a, b; Li and Wu, 2018; Zhang et al., 2020) and Transformer architectures (Vs et al., 2022; Ma et al., 2022; Tang et al., 2022; Xie et al., 2023; Tang et al., 2023a) for feature extraction and reconstruction. However, these models face limitations. However, both types of models have inherent limitations. Specifically, the utilization of static convolutional layers in these fusion methods results in uniform traversal of the entire image, limiting their ability to capture global contextual information due to constrained receptive fields (Ma et al., 2022). Moreover, the indiscriminate processing of each pixel by static convolutional layers overlooks nuanced spatial differences across various positions. Given the exigency of image fusion tasks necessitating diverse input data processing, current convolution-based techniques struggle to extract features effectively from different modalities, leading to diminished fusion performance. Transformer-based models excel in global modeling but suffer from quadratic complexity concerning image size due to their self-attention mechanism (Vaswani et al., 2017), imposing substantial computational overhead. Additionally, Transformers exhibit lower precision in capturing local features compared to CNNs. Although certain fusion models (Xie et al., 2023; Ma et al., 2022; Vs et al., 2022; Zhao et al., 2023) adopt a hybrid approach by amalgamating convolutional and Transformer layers to leverage their respective strengths and mitigate weaknesses, computational overhead remains a significant concern.
@@
Image fusion is an essential image enhancement technique (Ma et al., 2022). Methodologies often employ three main feature fusion strategies. Firstly, some methods use element-wise operations like addition, multiplication, or concatenation (Zhang et al., 2020; Li et al., 2020) to merge features from different modalities. However, these approaches overlook inter-modal relationships, leading to compromised fusion performance. Secondly, there are approaches that focus on deeper feature extraction (Vs et al., 2022; Xie et al., 2023), but they still lack effective inter-modal interaction and texture detail emphasis. Lastly, specific techniques utilize cross-modal fusion (Ma et al., 2022), including convolution-based and Transformer-based methods, which partially address feature interaction but have their limitations. Overall, contemporary fusion methodologies struggle to optimally utilize modal features and highlight crucial information, indicating a need for improved modal connections and emphasis on key image details.
@@
The advancement of Mamba (Gu and Dao, 2023) represents a promising avenue for achieving a balance between global receptive fields and computational efficiency. By formalizing the discretized state-space equations of Mamba into recursive form and incorporating specially designed structured reparameterization (Guo et al., 2024), it becomes capable of simulating very long-range dependencies. Moreover, the implementation of a parallel scanning algorithm (Ruan and Xiang, 2024) enables Mamba to process each token in a parallel manner, thereby facilitating efficient training on modern hardware like GPUs (Pei et al., 2024). These encouraging characteristics inspire us to further explore Mamba’s potential for efficient long-range modeling within image fusion networks.
@@
For the feature fusion part, we design a Dynamic Feature Fusion Module comprising a Dynamic Feature Enhancement Module and a Cross Modality Fusion Mamba Module. Unlike previous feature fusion methods, this fusion module dynamically enhances the detail texture information and differential information in source images and facilitates better information interaction between modalities. The Dynamic Feature Enhancement Module (DFEM), composed of dynamic differential convolution and dynamic difference-sensing attention, is used for adaptive feature enhancement. The DFEM module dynamically enhances critical information by establishing connections between input features from different modalities. We design a Cross Modality Fusion Mamba Module to effectively mine correlation information between modalities. Our network architecture adopts a Unet (Ronneberger et al., 2015) multi-layer structure, realizing an efficient and universal image fusion framework. Experimental results demonstrate that our proposed method outperforms state-of-the-art (SOTA) image fusion methods across various evaluation metrics on multiple multimodal baseline datasets, including IR-VIS fusion, CT-MRI fusion, PET-MRI fusion, SPECT-MRI fusion, and GFP-PC fusion.
@@ 2.1. Deep Multimodal Image Fusion
Image fusion, as a critical image enhancement technology, has garnered significant attention in recent years (Ma et al., 2022). In the era of deep learning, the primary image fusion methodologies can be categorized into four groups: Convolutional Neural Network (CNN) models (Zhang et al., 2020; Xu et al., 2020b, a), Generative Adversarial Network (GAN) models (Ma et al., 2019, 2020), Autoencoder (AE) based models (Li and Wu, 2018; Li et al., 2020, 2021; Ram Prabhakar et al., 2017), and models based on Transformer architectures (Xie et al., 2023; Vs et al., 2022; Ma et al., 2022; Tang et al., 2022; Zhao et al., 2023). These methods often employ straightforward fusion rules like element addition, averaging, or multiplication (Zhang et al., 2020). While these convolutional methods can achieve satisfactory fusion results, they are limited in their ability to extract global features due to inherent convolution constraints (Vs et al., 2022). Ma et al. (Ma et al., 2019) introduced an end-to-end GAN model named FusionGAN, where the source image is fed into the generation network, and the fusion result is generated by the discriminator network. Subsequently, they proposed a dual discriminator conditional generative adversarial network (DDcGAN) (Ma et al., 2020), which employs two discriminators to discern structural differences between the fused image and the two source images. While GANs demonstrate strong performance, their training process can be unstable, leading to potential texture distortion and difficulties in achieving stable fusion results due to the need to balance multi-source images using dual discriminators. Models based on autoencoders (AE) require a fusion rule to integrate features extracted from the encoder and generate the fusion result from the decoder. For example, RFN-Nest (Li et al., 2021) introduced a two-stage learning strategy and a customizable fusion approach. Furthermore, researchers have explored Transformer-based methods, with IFT (Vs et al., 2022) and CDDFuse (Zhao et al., 2023) effectively merging CNNs with Transformer architectures. SwinFusion (Ma et al., 2022) leverages SwinTransformer’s unique sliding window attention mechanism to address limitations in the convolutional fusion approach.
@@ 2.2. Mamba
State Space Models (SSMs) (Gu et al., 2021) have become a competitive backbone in deep learning, originating from classic control theory and offering linear scalability with sequence length for long-range dependency modeling. Structured State Space Sequence Models (S4) and Mamba (Liu et al., 2024), both rely on a classical continuous system that maps a 1D input function or sequence, denoted as $x(t)\in\mathbb{R}^{L}$, through intermediate implicit states $h(t)\in\mathbb{R}^{L}$ to an output $y(t)\in\mathbb{R}^{L}$. SSMs can be represented as the following linear Ordinary Differential Equation (ODE):
@@
Mamba has significantly advanced natural language tasks, surpassing traditional Transformers, with its data-dependent mechanisms, efficient hardware, and superior language processing. Expanding beyond language tasks, Mamba has also been applied successfully to vision tasks like image classification, video understanding, and biomedical image segmentation. This success has spurred a wave of research focusing on adapting Mamba-based models for specialized vision applications, including medical image segmentation with adaptations like Vm-unet (Ruan and Xiang, 2024). Additionally, Mamba has been integrated into graph representation tasks, enhancing graph embeddings and processing capabilities through models. Mamba’s versatility and efficiency make it a compelling choice for a wide range of applications, from language processing to computer vision representation tasks (Guo et al., 2024; Pei et al., 2024).
@@ 3.1. Motivation
The motivation of this paper lies in addressing the common problems and challenges in the field of multimodal image fusion. In recent years, with the recognition of the powerful capabilities of deep learning methods in image feature extraction, many fusion algorithms have emerged. However, the current mainstream convolutional neural networks (CNNs) (Zhang et al., 2020; Xu et al., 2020a, b) and Transformers models (Xie et al., 2023; Vs et al., 2022; Ma et al., 2022) have inherent limitations in image fusion, prompting us to delve into and design a new Dynamic Feature Enhancement Model to tackle these challenges and issues.
On the one hand, current fusion methods mainly use static convolutional layers for feature extraction and reconstruction, which have limitations in capturing nuanced spatial differences and global contextual information (Guan et al., 2023). On the other hand, Transformer models excel in global modeling, but their quadratic complexity in image size due to the self-attention mechanism restricts their computational efficiency (Ruan and Xiang, 2024). Additionally, they are less accurate in capturing local features compared to CNNs. Moreover, existing fusion methods also fall short in feature fusion, unable to effectively extract features from different modalities, leading to a decline in fusion performance.
@@ 3.3. Dynamic Vision State Space Module
We propose the Dynamic Vision State Space (DVSS) module as a modification of the SSM block for image fusion processing. In Fig. 2(b), starting with the input deep feature, we initially apply LayerNorm (LN) followed by the Efficient State Space Module (ESSM) (Pei et al., 2024) to capture spatial long-term dependencies.
Since SSMs process flattened feature maps as 1D token sequences, the chosen flattening strategy significantly impacts the number of neighboring pixels in the sequence. For example, when using the four-direction unfolding strategy, only four nearest neighbors are accessible to the anchor pixel. Specially, efficient 2D scan (ES2D, in Fig. 2(e)) (Pei et al., 2024) efficiently downscales 2D-Selective-Scan (SS2D) using skipping sampling and combines processed patches for global feature extraction. Consequently, some spatially close pixels in the 2D feature map become distant from each other in the 1D token sequence, potentially resulting in local pixel forgetting. To tackle this issue, we introduce an additional dynamic local convolution after ES2D (Pei et al., 2024) to restore neighborhood similarity. The Learnable Descriptive Convolution (LDC, in Fig. 2(f)) effectively learns intricate textural features, making it highly suitable in here. we employ LayerNorm for normalization and then utilize Learnable Descriptive Convolution (LDC) (Huang et al., 2022) to compensate for local features.
Moreover, SSMs often introduce a larger number of hidden states to remember long-range dependencies, leading to significant channel redundancy when visualizing the activation results for different channels (Liu et al., 2024). To enhance the expressive power of different channels, we integrate Efficient Channel Attention (ECA, in Fig. 2(g)) (Wang et al., 2020) into DVSS. This allows SSMs to concentrate on learning diverse channel representations, with subsequent channel attention selecting critical channels to prevent redundancy. DVSS is formulated as:
@@ 3.4. Dynamic Feature Enhancement Module
Within the module, the Learnable Descriptive Convolution (LDC) (Huang et al., 2022) module enhances texture processing on input feature maps using learnable mask parameters and convolution operations. By adjusting the convolution kernel’s weights, it emphasizes texture information, enhancing the model’s texture feature perception. Conversely, the Dynamic Difference Perception Attention calculates difference weights between input feature maps and applies them to both fused and original features, amplifying feature differences. This method helps the model effectively capture subtle differences among input features, leading to improved resolution and perception across different features and contributing to the model’s enhanced fusion performance. Finally, $\mathrm{DFEM}_{1}$ combines the texture enhancement feature, difference enhancement feature, and low-frequency feature ($\textbf{F}_{1}^{n}$) with background information, outputting $\textbf{D}_{1}^{n}$.
@@ 3.5. Cross Modality Fusion Mamba Module
These hybrid enhanced features are then input into the Efficient Spatial Scanning 2D (ES2D) (15) layer to capture the spatial long-term dependency.
@@ 3.6. Loss Function
SSIM (Cvejic et al., 2005) can measure the degree of distortion and the degree of similarity between two images. We use SSIM loss to guarantee structure similarity between the fused image and source images. It can be defined as
@@ 4.1.1. Datasets
we choose the KAIST (Hwang et al., 2015) dataset to train our FusionMamba. In this dataset, 70000 pairs of infrared and visible images are chosen for training. These images are also converted to gray scale and resized to 256 × 256. We use Harvard Medical Dataset  [^1] http://www.med.harvard.edu/AANLIB/home.html as the training and test dataset independently. There are 166 CT-MRI image pairs, 329 PET-MRI image pairs and 539 SPECT-MRI image pairs. The size of images is 256×256. We augment image pairs to 30000 via image rotation, which can improve the model and help to combat potential overfitting. 21 pairs of test images are randomly chosen to evaluate the model. The GFP database1 released by John Innes Centre (Koroleva et al., 2005) is employed in our experiments. The database contains 148 pairs of preregistered GFP and PC images of Arabidopsis thealiana cells with the same size of 358 × 358 pixels. We also augment image pairs to 30000 via image rotation. 20 pairs of test images are randomly chosen to evaluate the model.